OpenStudy (anonymous):

Solve for K 2(J+K)=M

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

a) divide by 2 b) subtract \(J\)

OpenStudy (anonymous):

So is it K=j/-2? @misty1212

OpenStudy (misty1212):

lets go slow

OpenStudy (misty1212):

\[2(J+K)=M\] divide by 2, which just means write it \[J+K=\frac{M}{2}\]

OpenStudy (misty1212):

subtract \(J\) again, just write it \[K=\frac{M}{2}-J\]

OpenStudy (anonymous):

aw okay I forgot about the M Thank you very much @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]